There are 16 blue cards and 9 red cards in a box. If Jane selects a card at random 200 times, about how many times would you expect her to pick a red card?
Answer by stanbon(75887)

There are 16 blue cards and 9 red cards in a box. If Jane selects a card at random 200 times, about how many times would you expect her to pick a red card?
------
mean = expected value = n*p = 200*(9/25) = 72
-------------
Cheers,
Stan H.